OpenStudy (anonymous):

Which shows 52^2 - 48^2 being evaluated using the difference of perfect squares method? A. 52^2 - 48^2 = (2704 + 2304)(2704 - 2304) = 2,003,200 B. 52^2 - 48^2 = (52 + 48)(52 - 48) = (100)(4) = 400 C. 52^2 - 48^2 = 2704 - 2304 = 400 D. 52^2 - 48^2 = (52 - 48)^2 = (4)^2 = 16

hartnn (hartnn):

a^2-b^2 =(a+b)(a-b) plug in a = 52, b= 48

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Thanks man

hartnn (hartnn):

no

hartnn (hartnn):

difference of squares, a^2 -b^2 = (a+b)(a-b) 52^2-48^2 = (52+48)(52-48)

hartnn (hartnn):

B.